OpenStudy (openstudygirl2):

Complete the equations to solve 350 ÷ 7. 350÷7 = (35÷7) x _______ =_______ Please help wil fan and medal

OpenStudy (3mar):

May I help?

OpenStudy (shadowlegendx):

You don't have to ask 3mar xD

OpenStudy (3mar):

Yes, I know but I need the asker to pay attention for help!

OpenStudy (3mar):

@openstudygirl2 Need help?

OpenStudy (shadowlegendx):

@openstudygirl2 hello

OpenStudy (openstudygirl2):

yes

OpenStudy (3mar):

Can you retype your equation so I can guide you well?

OpenStudy (shadowlegendx):

\[350 \div 7 = (35 \div 7) \times ? = ?\]

OpenStudy (shadowlegendx):

\[(35 \div 7) = (\frac{ 35 }{ 7 })\]

OpenStudy (3mar):

So it is your turn, Cute Cat!

OpenStudy (openstudygirl2):

yes

OpenStudy (openstudygirl2):

BTW im not great with math

OpenStudy (shadowlegendx):

\[350 \div 7 = \frac{ 350 }{ 7 } = \frac{ 35 }{ 7 } \times \frac{ 10 }{ 1 } = \frac{ 35 }{ 7 } \times 10\]

OpenStudy (openstudygirl2):



OpenStudy (shadowlegendx):

\[(35 \div 7) = \frac{ 35 }{ 7 }\]

OpenStudy (openstudygirl2):

35÷7 = 5

OpenStudy (shadowlegendx):

\[350 \div 7 = (35 \div 7) \times ? = ?\]

OpenStudy (shadowlegendx):

What goes in the first question mark, and what goes in the second question mark?

OpenStudy (openstudygirl2):

brb got my break be back in 15 minutes